A sample of N=9 scores has a mean of M=20. One of the scores is changed and the new mean is found to be 22. If the score was originally X=7, what is its new value.
1
Expert's answer
2020-12-04T12:15:12-0500

We have

n = 9

"\mu_{old}=20"

"\mu_{new}=22"

Lets denote the changing score "x_{old}=7" , so we need to find "x_{new}"


"\mu=\frac{1}{n}\sum\limits_{i=1}^9 x_i"

"\mu_{old}=\frac{1}{n}(\sum\limits_{i=1}^8 x_i + x_{old})"

"\mu_{new}=\frac{1}{n}(\sum\limits_{i=1}^8 x_i + x_{new})"

Thus we have

"\\begin{cases} 20=\\frac{1}{9}(\\sum\\limits_{i=1}^8+7) \\\\ 22=\\frac{1}{9}(\\sum\\limits_{i=1}^8+x_{new}) \\end{cases} \\implies \\begin{cases} \\sum\\limits_{i=1}^8+7=180 \\\\ \\sum\\limits_{i=1}^8+ x_{new}=198 \\end{cases}\\implies x_{new}=25\n\u200b"


Answer: 25.